If z = f(x, y), where x = eu + e−v , y = e−u − ev , then find ∂(u,v)/∂(x,y) .


1
Expert's answer
2021-11-08T16:29:42-0500

Consider: "z=f(x,y)," where "x=e^u+e^{-v},y=e^{-u}-e^v"

Require to find: "\frac{\partial (u,v)}{\partial (x,y)}"


Recollect the following: "\frac{\partial (u,v)}{\partial (x,y)}=\frac{1}{\frac{\partial (x,y)}{\partial (u,v)}}" and "\frac{\partial (x,y)}{\partial (u,v)}=\begin{vmatrix}
 \frac{\partial x}{\partial u}&\frac{\partial x}{\partial v} \\ 
 \frac{\partial y}{\partial u}& \frac{\partial y}{\partial v}
\end{vmatrix}"


Now "x=e^u+e^{-v},y=e^{-u}-e^v\Rightarrow \frac{\partial (x,y)}{\partial (u,v)}=\begin{vmatrix}
e^u & -e^{-v}\\ 
 -e^{-u}& -e^v
\end{vmatrix}"

"\Rightarrow \frac{\partial (x,y)}{\partial (u,v)}=-e^{u+v}-e^{-u-v}"


Using the above, we have

"\frac{\partial (u,v)}{\partial (x,y)}=\frac{1}{\frac{\partial (x,y)}{\partial (u,v)}}\Rightarrow \frac{\partial (u,v)}{\partial (x,y)}=\frac{1}{-e^{u+v}-e^{-u-v}}"


Therefore,

"\frac{\partial (u,v)}{\partial (x,y)}=\frac{1}{-e^{u+v}-e^{-u-v}}"